\begin{thebibliography}{99}
\bibitem{goodrich} Goodrich, Christopher, and Allan C. Peterson. Discrete
fractional calculus. Berlin: Springer, 2015.

\bibitem{ferhan} At\i c\i , Ferhan M., and Paul W. Eloe. "Discrete
fractional calculus with the nabla operator." Electronic Journal of
Qualitative Theory of Differential Equations 2009.3 (2009): 1-12.

\bibitem{ferhan3} Acar, Nihan, and Ferhan M. At\i c\i . "Exponential
functions of discrete fractional calculus." Applicable Analysis and Discrete
Mathematics (2013): 343-353.

\bibitem{ferhan4} Atici, Ferhan, and Paul Eloe. "Initial value problems in
discrete fractional calculus." Proceedings of the American Mathematical
Society 137.3 (2009): 981-989.

\bibitem{ferhan5} Atici, Ferhan M., and PAUL W. Eloe. "Linear forward
fractional difference equations." Commun. Appl. Anal 19 (2015): 31-42.

\bibitem{anas} Anastassiou, George A. "Nabla discrete fractional calculus
and nabla inequalities." Mathematical and Computer Modelling 51.5 (2010):
562-571.

\bibitem{anas2} Anastassiou, George A. "Right nabla discrete fractional
calculus." Int. J. Difference Equ 6.2 (2011): 91-104.

\bibitem{bohner} Bohner, Martin, and Allan C. Peterson, eds. Advances in
dynamic equations on time scales. Springer Science \& Business Media, 2002.

\bibitem{thabet} Abdeljawad, Thabet. "Dual identities in fractional
difference calculus within Riemann." Advances in Difference Equations 2013.1
(2013): 36.

\bibitem{thabet2} Abdeljawad, Thabet, and Dumitru Baleanu. "Fractional
Differences and Integration by Parts." Journal of Computational Analysis \&
Applications 13.3 (2011).

\bibitem{thabet3} Abdeljawad, Thabet. "On Riemann and Caputo fractional
differences." Computers \& Mathematics with Applications 62.3 (2011):
1602-1611.

\bibitem{thabet4} Abdeljawad, Thabet, and Ferhan M. Atici. "On the
definitions of nabla fractional operators." Abstract and Applied Analysis.
Vol. 2012. Hindawi Publishing Corporation, 2012.

\bibitem{thabet5} Abdeljawad, Thabet. "On delta and nabla Caputo fractional
differences and dual identities." Discrete Dynamics in Nature and Society
2013 (2013).

\bibitem{dumitru} Baleanu, Dumitru, Shahram Rezapour, and Saeid Salehi. "On
some self-adjoint fractional finite difference equations." Journal of
Computational Analysis \& Applications 19.1 (2015).

\bibitem{miller} Miller, Kenneth S., and Bertram Ross. "Fractional
difference calculus." Proceedings of the international symposium on
univalent functions, fractional calculus and their applications. 1988.

\bibitem{gray} Gray, Henry L., and Nien Fan Zhang. "On a new definition of
the fractional difference." Mathematics of Computation 50.182 (1988):
513-529.

\bibitem{diaz} Diaz, J. B., and T. J. Osler. "Differences of fractional
order." Mathematics of Computation 28.125 (1974): 185-202.

\bibitem{almeida} Almeida, Ricardo, et al. "Variational methods for the
solution of fractional discrete/continuous Sturm--Liouville problems."
Journal of Mechanics of Materials and Structures 12.1 (2016): 3-21.

\bibitem{bourdin} Bourdin, Lo\"{\i}c, et al. "Variational integrator for
fractional Euler--Lagrange equations." Applied Numerical Mathematics 71
(2013): 14-23.

\bibitem{fransiz} Dubois, Fran\c{c}ois, Ana Cristina Galucio, and Nelly
Point. "Introduction \`{a} la d\'{e}rivation fractionnaire. Th\'{e}orie et
applications." Techniques de l'ing\'{e}nieur (2010).

\bibitem{jin} Cheng, Jin-Fa, and Yu-Ming Chu. "Fractional difference
equations with real variable." Abstract and Applied Analysis. Vol. 2012.
Hindawi Publishing Corporation, 2012.

\bibitem{mohan} Mohan, J. Jagan, and G. V. S. R. Deekshitulu. "Fractional
order difference equations." International Journal of Differential Equations
2012 (2012).

\bibitem{bas} Bas, Erdal, and Funda Metin. "Fractional singular
Sturm-Liouville operator for Coulomb potential." Advances in Difference
Equations 2013.1 (2013): 300.

\bibitem{bas2} Bas, Erdal. "Fundamental spectral theory of fractional
singular Sturm-Liouville operator." Journal of Function Spaces and
Applications 2013 (2013).

\bibitem{klimek} Klimek, Malgorzata, and Om Prakash Agrawal. "Fractional
Sturm--Liouville problem." Computers \& Mathematics with Applications 66.5
(2013): 795-812.

\bibitem{klimek2} Klimek, Malgorzata, and Om P. Agrawal. "On a regular
fractional Sturm-Liouville problem with derivatives of order in (0, 1)."
Carpathian Control Conference (ICCC), 2012 13th International. IEEE, 2012.

\bibitem{lizama} Lizama, Carlos, Fractional Difference Equations and
Applications,
http://www.iumpa.upv.es/wp-content/uploads/2016/05/Fractional-Diference-Equations-and-Applications-2014.pdf

\bibitem{caputo} Ahrendt, Kevin - Rolling, Tim - Dewolf, Lydia - Mazurowski,
Liam - Mitchell, Kelsey \& Veconi, Dominic, Initial and Boundary Value
Problems for the Caputo Fractional Self-Adjoint Difference Equations,
Enlightenment of Pure and Applied Mathematics, Volume 2 ( 2016 ) , Issue 1,
Pages 105-141.
\end{thebibliography}

\documentclass{article}
\usepackage{amsfonts}

\input{tcilatex}
\begin{document}

\begin{center}
\textbf{DISCRETE FRACTIONAL STURM-LIOUVILLE EQUATIONS}

\bigskip

Erdal BAS$^{a}$ and Ramazan OZARSLAN$^{a\ast }$

\bigskip

$^{a}$\textit{Firat University, Science Faculty, Department of Mathematics,
23119 Elazig/Turkey}

\bigskip

\textit{e-mail: erdalmat@yahoo.com, }$^{\ast }$\textit{%
ozarslanramazan@gmail.com}

\bigskip
\end{center}

\noindent \textbf{Abstract}. In this study, we define discrete fractional
Sturm-Liouville (DFSL) operators within Riemann-Liouville and Gr\"{u}%
nwald-Letnikov fractional operators with both delta and nabla operators. We
show self-adjointness of the DFSL operator for the first time and prove some
spectral properties, like orthogonality of distinct eigenfunctions, reality
of eigenvalues, paralelly in integer and fractional order differential
operator counterparts.

\bigskip

\noindent \textbf{Keywords: }Sturm-Liouville, discrete fractional,
self-adjointness, eigenvalue, eigenfunction.

\bigskip

\noindent \textbf{AMS Subject Classification}: 34B24, 39A70, 34A08.

\bigskip

\begin{center}
\textbf{1. Introduction}

\bigskip
\end{center}

Fractional calculus is rather attractive subject due to having wide-ranging
application areas of theoretical and applied sciences. Although there is a
large number of a worthwhile mathematical works on the fractional
differential calculus, there is no noteworthy parallel improvement of
fractional difference calculus up to lately. This improvement has shown that
discrete fractional calculus has certain unforeseen hardship.

Fractional sums and differences were obtained firstly in Diaz-Osler \cite%
{diaz}, Miller and Ross \cite{miller} and Gray and Zhang \cite{gray} and
they found discrete types of fractional integrals and derivatives. Later,
several authors started to touch upon discrete fractional calculus \cite%
{goodrich, dumitru, lizama, mohan, caputo}. Nevertheless, discrete
fractional calculus is a rather novel area. The first studies has been done
by Atici et al. $\left[ 2-5\right] $, Anastassiou \cite{anas, anas2},
Abdeljawad et al. $\left[ 9-13\right] $, and Cheng \cite{jin}, and so forth.

Self-adjoint operators have an important place in differential operators.
Especially, various spectral properties of Sturm-Liouville differential
operators, like orthogonality of distinct eigenfunctions, reality of
eigenvalues, could be analyzed by the help of the self-adjointness.
Self-adjointness of difference operators enabled to be analyzed integer
order Sturm-Liouville difference operators parallelly to differential
counterpart.

Self-adjointness of fractional Sturm-Liouville differential operators have
been shown by $\left[ 23-26\right] $. However, although self-adjointness of
DFSL operators has been mentioned, it has not been proved yet. In this
study, we define DFSL operators, which is defined differently from \cite%
{almeida}, and prove the self-adjointness for the first time. Hereupon, we
prove some spectral properties of the operator paralelly in integer and
fractional order differential counterparts. More recently, Almeida et al. 
\cite{almeida} proved the orthogonality of distinct eigenfunctions, reality
of eigenvalues of DFSL operator and we prove these properties by using
self-adjointness.

In this study, differently we analyze DFSL equation within Riemann-Liouville
and Gr\"{u}nwald-Letnikov fractional operators with both delta and nabla
operators. The aim of this paper is to contribute to the theory of DFSL
operator.

In this paper, we discuss DFSL equation in two different ways;

\noindent $i)$ (nabla right and left) Riemann-Liouville (\textbf{R-L)}
fractional operator,%
\[
L_{1}x\left( t\right) =\nabla _{a}^{\mu }\left( p\left( t\right) _{b}\nabla
^{\mu }x\left( t\right) \right) +q\left( t\right) x\left( t\right) =\lambda
r\left( t\right) x\left( t\right) , 
\]

\noindent $ii)$ (delta left and right) Gr\"{u}nwald-Letnikov (\textbf{G-L)}
fractional operator,%
\[
L_{2}x\left( t\right) =\Delta _{-}^{\mu }\left( p\left( t\right) \Delta
_{+}^{\mu }x\left( t\right) \right) +q\left( t\right) x\left( t\right)
=\lambda r\left( t\right) x\left( t\right) . 
\]

\begin{center}
\bigskip

\textbf{2. Preliminaries}

\bigskip
\end{center}

\noindent \textbf{Definition 1.} \cite{goodrich} Delta and nabla difference
operators are defined by respectively%
\begin{equation}
\Delta x\left( t\right) =x\left( t+1\right) -x\left( t\right) ,\qquad \nabla
x\left( t\right) =x\left( t\right) -x\left( t-1\right) .  \tag{1}  \label{1}
\end{equation}

\bigskip

\noindent \textbf{Definition 2. }\cite{bohner} Falling function is defined
by, $\alpha \in 
%TCIMACRO{\U{211d} }%
%BeginExpansion
\mathbb{R}
%EndExpansion
,$ 
\begin{equation}
t^{\underline{\alpha }}=\frac{\Gamma \left( \alpha +1\right) }{\Gamma \left(
\alpha +1-n\right) },  \tag{2}  \label{2}
\end{equation}%
where $\Gamma $ is the gamma function.

\bigskip

\noindent \textbf{Definition 3. }\cite{bohner} Rising function is defined
by, $\alpha \in 
%TCIMACRO{\U{211d} }%
%BeginExpansion
\mathbb{R}
%EndExpansion
,$ 
\begin{equation}
t^{\bar{\alpha}}=\frac{\Gamma \left( t+\alpha \right) }{\Gamma \left(
t\right) }.  \tag{3}  \label{3}
\end{equation}%
Gamma function must be well-defined in the above two definitions.

\bigskip

\noindent \textbf{Remark 1. }Delta and nabla operators have the following
properties%
\begin{eqnarray}
\Delta t^{\underline{\alpha }} &=&\alpha t^{\underline{\alpha -1}}, 
\TCItag{4}  \label{4} \\
\nabla t^{\bar{\alpha}} &=&\alpha t^{\overline{\alpha -1}}.  \nonumber
\end{eqnarray}

\bigskip

\noindent \textbf{Definition 4. }\cite{ferhan, thabet, miller} Fractional
sum operators are defined by,

\noindent $\left( i\right) $ The nabla left fractional sum of order $\mu >0$
is defined 
\begin{equation}
\nabla _{a}^{-\mu }x\left( t\right) =\frac{1}{\Gamma \left( \mu \right) }%
\sum_{s=a+1}^{t}\left( t-\rho \left( s\right) \right) ^{\overline{\mu -1}%
}x\left( s\right) ,\text{ }t\in 
%TCIMACRO{\U{2115} }%
%BeginExpansion
\mathbb{N}
%EndExpansion
_{a+1},  \tag{5}  \label{6}
\end{equation}%
\noindent $\left( ii\right) $ The nabla right fractional sum of order $\mu
>0 $ is defined 
\begin{equation}
_{b}\nabla ^{-\mu }x\left( t\right) =\frac{1}{\Gamma \left( \mu \right) }%
\sum_{s=t}^{b-1}\left( s-\rho \left( t\right) \right) ^{\overline{\mu -1}%
}x\left( s\right) ,\text{ }t\in \text{ }_{b-1}%
%TCIMACRO{\U{2115} }%
%BeginExpansion
\mathbb{N}
%EndExpansion
,  \tag{6}  \label{7}
\end{equation}%
where $\rho \left( t\right) =t-1$ is called backward jump operators, $%
%TCIMACRO{\U{2115} }%
%BeginExpansion
\mathbb{N}
%EndExpansion
_{a}=\left \{ a,a+1,...\right \} ,$ $_{b}%
%TCIMACRO{\U{2115} }%
%BeginExpansion
\mathbb{N}
%EndExpansion
=\left \{ b,b-1,...\right \} $.

\bigskip

\noindent \textbf{Definition 5. }\cite{thabet2, thabet4} Fractional
difference operators are defined by,

\noindent $\left( i\right) $ The nabla left fractional difference of order $%
\mu >0$ is defined 
\begin{equation}
\nabla _{a}^{\mu }x\left( t\right) =\nabla _{a}^{n}\nabla _{a}^{-\left(
n-\mu \right) }x\left( t\right) =\frac{\nabla ^{n}}{\Gamma \left( n-\mu
\right) }\sum_{s=a+1}^{t}\left( t-\rho \left( s\right) \right) ^{\overline{%
n-\mu -1}}x\left( s\right) ,\text{ }t\in 
%TCIMACRO{\U{2115} }%
%BeginExpansion
\mathbb{N}
%EndExpansion
_{a+1},  \tag{7}  \label{9}
\end{equation}%
$\noindent \left( ii\right) $ The nabla right fractional difference of order 
$\mu >0$ is defined 
\begin{equation}
_{b}\nabla ^{\mu }x\left( t\right) =\left( -1\right) ^{n}\nabla
_{b}^{n}\nabla _{a}^{-\left( n-\mu \right) }x\left( t\right) =\frac{\left(
-1\right) ^{n}\Delta ^{n}}{\Gamma \left( n-\mu \right) }\sum_{s=a+1}^{t}%
\left( s-\rho \left( t\right) \right) ^{\overline{n-\mu -1}}x\left( s\right)
,\text{ }t\in \text{ }_{b-1}%
%TCIMACRO{\U{2115} }%
%BeginExpansion
\mathbb{N}
%EndExpansion
.  \tag{8}  \label{10}
\end{equation}%
Fractional differences in $\left( \ref{9}-\ref{10}\right) $ are called the 
\textbf{Riemann--Liouville (R-L) }definition of the $\mu $-th order nabla
fractional difference.

\bigskip

\noindent \textbf{Definition 6. }$\left[ 18-20\right] $ Fractional
difference operators are defined by,

\noindent $\left( i\right) $ The delta left fractional difference of order $%
\mu ,$ $0<\mu \leq 1,$ is defined%
\begin{equation}
\Delta _{-}^{\mu }x\left( t\right) =\frac{1}{h^{\mu }}\sum_{s=0}^{t}\left(
-1\right) ^{s}\frac{\mu \left( \mu -1\right) ...\left( \mu -s+1\right) }{s!}%
x\left( t-s\right) ,\text{ }t=1,...,N.  \tag{9}  \label{11}
\end{equation}%
$\noindent \left( ii\right) $ The delta right fractional difference of order 
$\mu ,$ $0<\mu \leq 1,$ is defined%
\begin{equation}
\Delta _{+}^{\mu }x\left( t\right) =\frac{1}{h^{\mu }}\sum_{s=0}^{N-t}\left(
-1\right) ^{s}\frac{\mu \left( \mu -1\right) ...\left( \mu -s+1\right) }{s!}%
x\left( t+s\right) ,\text{ }t=0,..,N-1,  \tag{10}  \label{12}
\end{equation}%
Fractional differences in $\left( \ref{11}-\ref{12}\right) $ are called the 
\textbf{Gr\"{u}nwald--Letnikov (G-L) }definition of the $\mu $-th order
delta fractional difference.

\bigskip

\noindent \textbf{Definition 7. }\cite{thabet4} We define the integration by
parts formula for\textbf{\ R-L} nabla fractional difference operator, $u$ is
defined on $_{b}%
%TCIMACRO{\U{2115} }%
%BeginExpansion
\mathbb{N}
%EndExpansion
$ and $v$ is defined on $%
%TCIMACRO{\U{2115} }%
%BeginExpansion
\mathbb{N}
%EndExpansion
_{a}$, then 
\begin{equation}
\sum_{s=a+1}^{b-1}u\left( s\right) \nabla _{a}^{\mu }v\left( s\right)
=\sum_{s=a+1}^{b-1}v\left( s\right) _{b}\nabla ^{\mu }u\left( s\right) . 
\tag{11}  \label{16}
\end{equation}

\bigskip

\noindent \textbf{Definition 8. }\cite{almeida, bourdin} We define the
integration by parts formula for \textbf{G-L }delta fractional difference
operator, $u$, $v$ is defined on $\left \{ 0,1,...,n\right \} $, then 
\begin{equation}
\sum_{s=0}^{n}u\left( s\right) \Delta _{-}^{\mu }v\left( s\right)
=\sum_{s=0}^{n}v\left( s\right) \Delta _{+}^{\mu }u\left( s\right) . 
\tag{12}  \label{17}
\end{equation}

\begin{center}
\bigskip

\textbf{3. Main Results}
\end{center}

\bigskip

\noindent \textbf{3.1. Discrete Fractional Sturm-Liouville Equations}

\bigskip

We consider DFSL equations in two different ways;

\noindent $i)$ (nabla right and left) \textbf{R-L} fractional operator is
defined by,%
\begin{equation}
L_{1}x\left( t\right) =\nabla _{a}^{\mu }\left( p\left( t\right) _{b}\nabla
^{\mu }x\left( t\right) \right) +q\left( t\right) x\left( t\right) =\lambda
r\left( t\right) x\left( t\right) ,  \tag{13}  \label{23}
\end{equation}%
where $p\left( t\right) >0,$ $r\left( t\right) >0,$ $q\left( t\right) $ is
defined and real valued, $\lambda $ is the spectral parameter, $x\left(
t\right) \in l^{2}\left[ a+1,b-1\right] .$

\noindent $ii)$ (delta left and right) \textbf{G-L} fractional operator is
defined by,%
\begin{equation}
L_{2}x\left( t\right) =\Delta _{-}^{\mu }\left( p\left( t\right) \Delta
_{+}^{\mu }x\left( t\right) \right) +q\left( t\right) x\left( t\right)
=\lambda r\left( t\right) x\left( t\right) ,  \tag{14}  \label{27}
\end{equation}%
where $p,q,r,\lambda $ is as defined above, $x\left( t\right) \in l^{2}\left[
0,n\right] .$

\bigskip

Firstly, let's consider the equation $\left( \ref{23}\right) $ and give the
following theorems and proofs$,$

\bigskip

\noindent \textbf{Theorem 9. }DFSL operator $L_{1}$, denoted by the equation 
$\left( \ref{23}\right) ,$ is self-adjoint.

\bigskip

\textbf{Proof.}%
\begin{eqnarray}
u\left( t\right) L_{1}v\left( t\right) &\mathbf{=}&u\left( t\right) \nabla
_{a}^{\mu }\left( p\left( t\right) _{b}\nabla ^{\mu }v\left( t\right)
\right) +u\left( t\right) q\left( t\right) v\left( t\right) ,  \TCItag{15} \\
v\left( t\right) L_{1}u\left( t\right) &\mathbf{=}&v\left( t\right) \nabla
_{a}^{\mu }\left( p\left( t\right) _{b}\nabla ^{\mu }u\left( t\right)
\right) +v\left( t\right) q\left( t\right) u\left( t\right) .  \TCItag{16}
\end{eqnarray}%
If $\left( 15-16\right) $ is subtracted from each other%
\[
u\left( t\right) L_{1}v\left( t\right) -v\left( t\right) L_{1}u\left(
t\right) =u\left( t\right) \nabla _{a}^{\mu }\left( p\left( t\right)
_{b}\nabla ^{\mu }v\left( t\right) \right) -v\left( t\right) \nabla
_{a}^{\mu }\left( p\left( t\right) _{b}\nabla ^{\mu }u\left( t\right)
\right) 
\]%
and definite sum operator to the both side of the last equality is applied,
we have%
\begin{equation}
\sum_{s=a+1}^{b-1}\left( u\left( s\right) L_{1}v\left( s\right) -v\left(
s\right) L_{1}u\left( s\right) \right) =\sum_{s=a+1}^{b-1}u\left( s\right)
\nabla _{a}^{\mu }\left( p\left( s\right) _{b}\nabla ^{\mu }v\left( s\right)
\right) -\sum_{s=a+1}^{b-1}v\left( s\right) \nabla _{a}^{\mu }\left( p\left(
s\right) _{b}\nabla ^{\mu }u\left( s\right) \right) .  \tag{17}
\end{equation}%
If we apply the integration by parts formula in $\left( \ref{16}\right) $ to
right hand side of $\left( 17\right) ,$ we have%
\begin{eqnarray*}
\sum_{s=a+1}^{b-1}\left( u\left( s\right) L_{1}v\left( s\right) -v\left(
s\right) L_{1}u\left( s\right) \right) &=&\sum_{s=a+1}^{b-1}p\left( s\right)
_{b}\nabla ^{\mu }v\left( s\right) _{b}\nabla ^{\mu }u\left( s\right) \\
&&-\sum_{s=a+1}^{b-1}p\left( s\right) _{b}\nabla ^{\mu }u\left( s\right)
_{b}\nabla ^{\mu }v\left( s\right) \\
&=&0,
\end{eqnarray*}%
\[
\left \langle L_{1}u,v\right \rangle =\left \langle u,L_{1}v\right \rangle . 
\]%
The proof completes.

\bigskip

\noindent \textbf{Theorem 10. }Eigenfunctions, corresponding to distinct
eigenvalues, of the equation $\left( \ref{27}\right) $ are orthogonal.

\bigskip

\textbf{Proof. }Let $\lambda _{\alpha }$ and $\lambda _{\beta }$ are two
different eigenvalues corresponds to eigenfunctions $u\left( n\right) $ and $%
v\left( n\right) $ respectively for the the equation $\left( \ref{23}\right) 
$,%
\begin{eqnarray*}
\nabla _{a}^{\mu }\left( p\left( t\right) _{b}\nabla ^{\mu }u\left( t\right)
\right) +q\left( t\right) u\left( t\right) -\lambda _{\alpha }r\left(
t\right) u\left( t\right) &=&0, \\
\nabla _{a}^{\mu }\left( p\left( t\right) _{b}\nabla ^{\mu }v\left( t\right)
\right) +q\left( t\right) v\left( t\right) -\lambda _{\beta }r\left(
t\right) v\left( t\right) &=&0,
\end{eqnarray*}%
If we multiply last two equations to $v\left( n\right) $ and $u\left(
n\right) $ respectively, subtract from each other and apply definite sum
operator, since the self-adjointness of the operator $L_{1},$ we get 
\[
\left( \lambda _{\alpha }-\lambda _{\beta }\right) \sum_{s=a+1}^{b-1}r\left(
s\right) u\left( s\right) v\left( s\right) =0, 
\]%
since $\lambda _{\alpha }\neq \lambda _{\beta ,}$%
\begin{eqnarray*}
\sum_{s=a+1}^{b-1}r\left( s\right) u\left( s\right) v\left( s\right) &=&0 \\
\left \langle u\left( t\right) ,v\left( t\right) \right \rangle &=&0
\end{eqnarray*}%
The proof completes..

\bigskip

\noindent \textbf{Theorem 11. }All eigenvalues of the equation $\left( \ref%
{23}\right) $ are real.

\bigskip

\textbf{Proof. }Let $\lambda =\alpha +i\beta ,$ since the self-adjointness
of the operator $L_{1},$ we have%
\begin{eqnarray*}
\left \langle L_{1}u,u\right \rangle &=&\left \langle u,L_{1}u\right \rangle
, \\
\left \langle \lambda ru,u\right \rangle &=&\left \langle u,\lambda ru\right
\rangle ,
\end{eqnarray*}%
\[
\left( \lambda -\overline{\lambda }\right) \left \langle u,u\right \rangle
_{r}=0 
\]%
Since $\left \langle u,u\right \rangle _{r}\neq 0,$ 
\[
\lambda =\overline{\lambda } 
\]%
and hence $\beta =0.$ The proof completes.

\bigskip

Secondly, let's consider the equation $\left( \ref{23}\right) $ and give the
following theorems and proofs$,$

\bigskip

\noindent \textbf{Theorem 12. }DFSL operator $L_{2}$, denoted by the
equation $\left( \ref{27}\right) ,$ is self-adjoint.

\bigskip

\textbf{Proof. }%
\begin{eqnarray}
u\left( t\right) L_{2}v\left( t\right) &\mathbf{=}&u\left( t\right) \Delta
_{-}^{\mu }\left( p\left( t\right) \Delta _{+}^{\mu }v\left( t\right)
\right) +u\left( t\right) q\left( t\right) v\left( t\right) ,  \TCItag{18} \\
v\left( t\right) L_{2}u\left( t\right) &\mathbf{=}&v\left( t\right) \Delta
_{-}^{\mu }\left( p\left( t\right) \Delta _{+}^{\mu }u\left( t\right)
\right) +v\left( t\right) q\left( t\right) u\left( t\right) .  \TCItag{19}
\end{eqnarray}%
If $\left( 18-19\right) $ is subtracted from each other%
\[
u\left( t\right) L_{2}v\left( t\right) -v\left( t\right) L_{2}u\left(
t\right) =u\left( t\right) \Delta _{-}^{\mu }\left( p\left( t\right) \Delta
_{+}^{\mu }v\left( t\right) \right) -v\left( t\right) \Delta _{-}^{\mu
}\left( p\left( t\right) \Delta _{+}^{\mu }u\left( t\right) \right) 
\]%
and definite sum operator to the both side of the last equality is applied,
we have%
\begin{equation}
\sum_{s=0}^{n}\left( u\left( s\right) L_{1}v\left( s\right) -v\left(
s\right) L_{2}u\left( s\right) \right) =\sum_{s=0}^{n}u\left( s\right)
\Delta _{-}^{\mu }\left( p\left( s\right) \Delta _{+}^{\mu }v\left( s\right)
\right) -\sum_{s=0}^{n}v\left( s\right) \Delta _{-}^{\mu }\left( p\left(
s\right) \Delta _{+}^{\mu }u\left( s\right) \right) .  \tag{20}
\end{equation}%
If we apply the integration by parts formula in $\left( \ref{17}\right) $ to
right hand side of $\left( 20\right) ,$ we have%
\begin{eqnarray*}
\sum_{s=0}^{n}\left( u\left( s\right) L_{2}v\left( s\right) -v\left(
s\right) L_{2}u\left( s\right) \right) &=&\sum_{s=0}^{n}p\left( s\right)
\Delta _{+}^{\mu }v\left( s\right) \Delta _{+}^{\mu }u\left( s\right) \\
&&-\sum_{s=0}^{n}p\left( s\right) \Delta _{+}^{\mu }u\left( s\right) \Delta
_{+}^{\mu }v\left( s\right) \\
&=&0
\end{eqnarray*}%
\[
\left \langle L_{2}u,v\right \rangle =\left \langle u,L_{2}v\right \rangle . 
\]%
The proof completes.

\bigskip

\noindent \textbf{Theorem 13. }Eigenfunctions, corresponding to distinct
eigenvalues, of the equation $\left( \ref{27}\right) $ are orthogonal.

\bigskip

\textbf{Proof. }Let $\lambda _{\alpha }$ and $\lambda _{\beta }$ are two
different eigenvalues corresponds to eigenfunctions $u\left( n\right) $ and $%
v\left( n\right) $ respectively for the the equation $\left( \ref{27}\right) 
$,%
\begin{eqnarray*}
\Delta _{-}^{\mu }\left( p\left( t\right) \Delta _{+}^{\mu }u\left( t\right)
\right) +q\left( t\right) u\left( t\right) -\lambda _{\alpha }r\left(
t\right) u\left( t\right) &=&0, \\
\Delta _{-}^{\mu }\left( p\left( t\right) \Delta _{+}^{\mu }v\left( t\right)
\right) +q\left( t\right) v\left( t\right) -\lambda _{\beta }r\left(
t\right) v\left( t\right) &=&0.
\end{eqnarray*}%
If we multiply last two equations to $v\left( n\right) $ and $u\left(
n\right) $ respectively, subtract from each other and apply definite sum
operator, since the self-adjointness of the operator $L_{2},$ we get 
\[
\left( \lambda _{\alpha }-\lambda _{\beta }\right) \sum_{s=0}^{n}r\left(
s\right) u\left( s\right) v\left( s\right) =0, 
\]%
since $\lambda _{\alpha }\neq \lambda _{\beta ,}$%
\begin{eqnarray*}
\sum_{s=0}^{n}r\left( s\right) u\left( s\right) v\left( s\right) &=&0 \\
\left \langle u\left( t\right) ,v\left( t\right) \right \rangle &=&0.
\end{eqnarray*}%
So, the eigenfunctions are orthogonal. The proof completes.

\bigskip

\noindent \textbf{Theorem 14. }All eigenvalues of the equation $\left( \ref%
{27}\right) $ are real.

\bigskip

\textbf{Proof }Let $\lambda =\alpha +i\beta ,$ since the self-adjointness of
the operator $L_{2}$%
\begin{eqnarray*}
\left \langle L_{2}u,u\right \rangle &=&\left \langle u,L_{2}u\right \rangle
, \\
\left \langle \lambda ru,u\right \rangle &=&\left \langle u,\lambda ru\right
\rangle ,
\end{eqnarray*}%
\[
\left( \lambda -\overline{\lambda }\right) \left \langle u,u\right \rangle
_{r}=0 
\]%
Since $\left \langle u,u\right \rangle _{r}\neq 0,$ 
\[
\lambda =\overline{\lambda }, 
\]%
and hence $\beta =0.$ The proof completes.

\bigskip

\begin{center}
\textbf{4. Conclusion}
\end{center}

In this study, differently we analyze DFSL equation within Riemann-Liouville
and Gr\"{u}nwald-Letnikov fractional operators with both delta and nabla
operators. We define DFSL operators, which is defined differently from \cite%
{almeida}, and prove the self-adjointness for the first time. Hereupon, we
prove some spectral properties of the operator paralelly in integer and
fractional order differential counterparts, like orthogonality of distinct
eigenfunctions, reality of eigenvalues. The aim of this paper is to
contribute to the theory of Sturm-Liouville fractional difference operator.

\end{document}